# Some Robustness Properties of Label Cleaning

### v1

@@ 1 Introduction
Important contributions to the theory of surrogate risk consistency trace to the 2000s [46, 22, 34], with Bartlett, Jordan, and McAuliffe [3] characterizing when fitting a model using a convex surrogate is consistent for binary classification for the zero-one error. Since this work, there has been an abundance of work on surrogate risk consistency, including on multi-label classification [44, 37, 14, 43, 1], ranking problems [10, 11, 30], structured prediction [28, 4, 26], ordinal regression [29], and general theory [34]. On the one hand, these analyses, which consider the standard supervised learning scenario of data pairs $(X,Y)$, enable us to fully exploit the entire statistical theory of empirical processes [39, 2, 19, 3]. On the other, they do not address the data aggregation machinery now common in modern dataset creation.
@@ 2.1 Label aggregation
Instead of obtaining $(X_{i},Y_{i})$ pairs, consider the case that we replace the output $Y$ with a more abstract variable $Z\in\mathcal{Z}$. For example, in the motivating scenario in the introduction in which we collect multiple (say, $m$) noisy labels for each example $X$, we take $Z=(Y_{1},\ldots,Y_{m})\in\mathcal{Y}^{m}$. For an abstract “aggregation space” $\mathcal{A}$, let $A:\mathcal{Z}\to\mathcal{A}$ be an aggregating function (e.g., majority vote), and let $\varphi:\mathbb{R}^{d}\times\mathcal{A}\to\mathbb{R}_{+}$ be a surrogate loss defined on this aggregation space. We then define the aggregated surrogate risk
@@ 3 Surrogate consistency
Consistency and comparison inequalities follow from the calibration functions (see [46, Prop. 25] and [34, Thm. 2.8 and Lemma 2.9]):
@@ 3.1 Fisher consistency failure without label aggregation: ranking
Consider the problem of ranking $k$ items using pairwise comparison data [18, 13, 11, 24], where the space $\mathcal{Y}$ consists of all pairwise comparisons of these items, $\mathcal{Y}=\{(i,j):i\neq j,1\leq i,j\leq k\}$. The (population) rank aggregation problem is, for each $x$, to transform the probabilities $p_{ij}=P(Y=(i,j)\mid x)$ into a ranking of the $k$ items. While numerous possibilities exist for such aggregation, we consider a simple comparison-based aggregation scheme [cf. 18]; similar negative results to the one we show below hold for more sophisticated schemes. Define the normalized transition matrix $C_{x}\in\mathbb{R}_{+}^{k\times k}$ with entries $(C_{x})_{ii}=0$ and
@@ 3.2 Label aggregation obtains stronger surrogate consistency
The extended ranking example in ranking suggests potential benefits of aggregating labels, and it is natural to ask how aggregation interacts with surrogate consistency more generally. Thus, we present two results here: one that performs an essentially basic extension of standard surrogate-risk consistency, and the second that shows how aggregation-based methods can “upgrade” what might nominally be inconsistent losses into consistent losses, as Proposition 2 suggests may be possbile.
@@ 3.2.1 Basic extensions of surrogate consistency
A consistency result then follows, similar to Corollary 3.1, under appropriate measurability conditions (we will leave these tacit as they are not central to our results). Then more or less as a corollary of Steinwart [34, Thm. 2.8], we have the following consistency result. (We include a proof for completeness in Appendix B.1.)
@@ 3.2.3 Identifying surrogates and consistency amplification
Here, $\alpha\in[0,1]$ and $\beta\in[0,\infty]$, so that conditions (Nα) and (Mβ) always trivially hold with $\alpha=\beta=0$, moreover, as in the binary case [3, Thm. 3], they are equivalent via the transformation $\beta=\frac{\alpha}{1-\alpha}$. (See Appendix B.2.) We shall also use a noise condition number
@@
To provide a somewhat more concrete bound, we optimize over $M$ in Theorem 1, using the crued bound $\kappa(x)\leq 1/\Delta(x)$ on the condition number. By taking $M=1$ for $\mathop{\rm card}(\mathcal{Y})=k=2$ and optimizing $M$ for $k\geq 3$, we may lower bound $\xi_{m,k}$ in the level $(\xi_{m,k},\zeta)$-consistency (Def. 3.2) that in Theorem 1 promises, setting
@@ 3.3 Surrogate consistency examples with majority vote
We collect several examples, of varying levels of concreteness, that allow us to instatiate Theorem 1 and Corollary 3.3. Throughout, we shall assume that $P$ has a noise exponent $\alpha\in[0,1]$, though this is no loss of generality, as Condition (Nα) always holds with $\alpha=0$. We defer proofs for each result in this section to Appendix B.5.
@@ 4 Robustness and consistency for models
The previous section builds off of the now classical theory of surrogate risk consistency, which assumes $\mathcal{F}$ to be the class of all measurable functions. The results there show that aggregation can allow us to “upgrade” consistency so that even if a surrogate $\varphi$ is inconsistent for paired (non-aggregated) data $(X,Y)$, we can achieve level-$(\xi,\zeta)$ consistency (Def. 3.2) with sufficient aggregation. Here, we take a different view of the problem of consistency, considering the consequences of optimizing over a restricted (often parametric) hypothesis class $\mathcal{F}$. Of course, in a well-specified model, obtaining consistency with such a restricted hypothesis class is no issue, but it is unrealistic to assume such a brittle condition. This gives rise to the long-standing challenge of quantifying surrogate consistency when the hypothesis class contains only a subset of the measurable functions [12, 25]. We tackle some of the issues around this, showing that aggregating labels allows consistent estimates in scenarios where consistency might otherwise fail.
@@ 4.1 Consistency failure for binary classification in finite dimensions
Data aggregation methods provide one way to circumvent the the inconsistency Proposition 4 highlights. To state the result, define the approximate minimizers
@@ 4.2 Aggregation, consistency, and restricted hypothesis classes
for $\phi$ convex, non-increasing with $\phi^{\prime}(0)<0$, and $\inf_{t}\phi(t)=0$. Zhang [45, Thm. 5] shows that any such loss is consistent over the class $\mathcal{F}=\{f:\mathcal{X}\to\mathbb{R}^{k}\mid\mathbf{1}^{T}f=0\}$. Clearly, the margin-based binary setting in Sec. 4.1 falls into this scenario when we take f(x)=(g(x),g(x))f(x)=(g(x),-g(x)) for a measurable $g$. Additionally, in a parametric setting when $\mathcal{X}=\mathbb{R}^{d}$, if $\mathcal{F}$ consists of linear functions $f(x)=(\langle\theta_{1},x\rangle,\ldots,\langle\theta_{k},x\rangle)$ with $\sum_{i=1}^{k}\theta_{i}=0$, then $\mathcal{F}$ is a (convex) cone.
@@ 4.3 On finite-dimensional multiclass classification
$\varphi(\Theta^{\top}x,y)=-\langle\theta_{y},x\rangle+\log\left({1+\sum_{i=1}^{k-1}\exp(\langle\theta_{i},x\rangle)}\right),$
@@
Standard results in statistics guarantees both consistency and efficiency when the model is well-specified without aggregation, and when $m\geq 1$ and $k=2$, Cheng et al. [5, Prop. 3] show there exists $t_{m}>0$ such that $\Theta_{m}=t_{m}\Theta^{\star}$ if $X\sim\mathsf{N}(0,I_{d})$ even with a mis-specified link. This implies that in binary classification, even if the link function is incorrect, we can still achieve consistent classification regardless of the aggregation level $m$, as the direction $\Theta^{\star}/\left\|{\Theta^{\star}}\right\|_{2}$ determines consistency. However, as soon as $k\geq 3$ and the true link is slightly mis-specified, risk consistency fails. Fixing a set $\mathcal{T}_{\epsilon}\subset\mathbb{R}^{k-1}$ with Lebesgue measure $\epsilon$, consider
@@ A.1 Proof of Proposition 1
Proof  First, we observe that $g_{n}\to g$ pointwise implies that $g_{n}\to g$ uniformly on compacta, and $g$ is convex (see Hiriart-Urruty and Lemaréchal [16, Thm. IV.3.15] and Rockafellar and Wets [32, Thm. 7.17]). This is then equivalent to epigraphical convergence of $g_{n}$ to $g$ [32, Thm. 7.17]. Moreover, as $g_{n}\to g$ uniformly on compacta, if $x_{n}\to x$ then $g_{n}(x_{n})\to g(x)$. Thus, for any $\epsilon_{n}\downarrow 0$, if for a subsequence $n(m)\subset\mathbb{N}$ we have $x_{n(m)}\in\mathop{\epsilon_{n(m)}\textup{-argmin}}g_{n}$, and $x_{n(m)}\to x$, we certainly have $g_{n(m)}(x_{n(m)})\to g(x)$. Consequently [32, Prop. 7.30] we have
@@
That the limit supremum is non-empty is then a consequence of [32, Thm. 7.33], as the convex functions $g_{n}$ must be coercive as they are convex and $g$ is. ∎
@@ B.2 The equivalence of the Mammen-Tsybakov conditions (Nα) and (Mβ)
We show the analogue of Bartlett et al. [3, Thm. 3], essentially mimicking their proof, but including it for completeness.
@@ B.3 Proof of Theorem 1
Consider the second term in the bound (20). For any convex function $0\leq\psi$ with $\psi(0)=0$, $\epsilon\mapsto\psi(\epsilon)/\epsilon$ is non-decreasing on $\epsilon>0$ (cf. [16, Ch. 1]). This implies
@@ B.5 Proofs for the identifiable surrogate losses
That $\phi(\delta)<0$ is immediate because $\phi$ is non-increasing by assumption, and the monotonicity properties of convex functions [16, Ch. 1] guarantee it strictly decreases near 0. For $y\in\{\pm 1\}$, we take $s_{y}=\delta y$ and $a_{y}=y$, and
@@ D.2 Proof of Theorem 2
the convex conjugate of the gap functional on well-separated distributions. Then Zhang [45, Prop. 25] shows that $\overline{\psi}_{\varphi}^{**}(\epsilon)>0$ whenever $\overline{\psi}_{\varphi}(\epsilon)>0$.

### v2

@@ 1 Introduction
Important contributions to the theory of surrogate risk consistency trace to the 2000s [52, 24, 40], with Bartlett, Jordan, and McAuliffe [5] characterizing when fitting a model using a convex surrogate is consistent for binary classification for the zero-one error. Since this work, there has been an abundance of work on surrogate risk consistency, including on multi-label classification [50, 43, 16, 49, 1], ranking problems [12, 13, 36], structured prediction [34, 6, 32], ordinal regression [35], restricted hypothesis class [2, 3, 27, 29] and general theory [40]. On the one hand, these analyses, which consider the standard supervised learning scenario of data pairs $(X,Y)$, enable us to fully exploit the entire statistical theory of empirical processes [45, 4, 21, 5]. On the other, they do not address the data aggregation machinery now common in modern dataset creation.
@@ 2.1 Label aggregation
Instead of obtaining $(X_{i},Y_{i})$ pairs, consider the case that we replace the output $Y$ with a more abstract variable $Z\in\mathcal{Z}$. For example, in the motivating scenario in the introduction in which we collect multiple (say, $m$) noisy labels for each example $X$, we take $Z=(Y_{1},\ldots,Y_{m})\in\mathcal{Y}^{m}$. Mao et al. [27] consider similar multilabeled data structure for multilabel losses, while we focus on the label aggregation pipeline. For an abstract “aggregation space” $\mathcal{A}$, let $A:\mathcal{Z}\to\mathcal{A}$ be an aggregating function (e.g., majority vote), and let $\varphi:\mathbb{R}^{d}\times\mathcal{A}\to\mathbb{R}_{+}$ be a surrogate loss defined on this aggregation space. We then define the aggregated surrogate risk
@@ 3 Surrogate consistency
Consistency and comparison inequalities follow from the calibration functions (see [52, Prop. 25] and [40, Thm. 2.8 and Lemma 2.9]):
@@ 3.1 Fisher consistency failure without label aggregation: ranking
Consider the problem of ranking $k$ items using pairwise comparison data [20, 15, 13, 30], where the space $\mathcal{Y}$ consists of all pairwise comparisons of these items, $\mathcal{Y}=\{(i,j):i\neq j,1\leq i,j\leq k\}$. The (population) rank aggregation problem is, for each $x$, to transform the probabilities $p_{ij}=P(Y=(i,j)\mid x)$ into a ranking of the $k$ items. While numerous possibilities exist for such aggregation, we consider a simple comparison-based aggregation scheme [cf. 20]; similar negative results to the one we show below hold for more sophisticated schemes. Define the normalized transition matrix $C_{x}\in\mathbb{R}_{+}^{k\times k}$ with entries $(C_{x})_{ii}=0$ and
@@ 3.2 Label aggregation obtains stronger surrogate consistency
The extended ranking example in ranking suggests potential benefits of aggregating labels, and it is natural to ask how aggregation interacts with surrogate consistency more generally. Thus, we present two results here: one that performs an essentially basic extension of standard surrogate-risk consistency, and the second that shows how aggregation-based methods can “upgrade” what might nominally be inconsistent losses into consistent losses, as Proposition 2 suggests may be possible.
@@ 3.2.1 Basic extensions of surrogate consistency
A consistency result then follows, similar to Corollary 3.1, under appropriate measurability conditions (we will leave these tacit as they are not central to our results). Then more or less as a corollary of Steinwart [40, Thm. 2.8], we have the following consistency result. (We include a proof for completeness in Appendix B.1.)
@@ 3.2.3 Identifying surrogates and consistency amplification
Here, $\alpha\in[0,1]$ and $\beta\in[0,\infty]$, so that conditions (Nα) and (Mβ) always trivially hold with $\alpha=\beta=0$, moreover, as in the binary case [5, Thm. 3], they are equivalent via the transformation $\beta=\frac{\alpha}{1-\alpha}$. (See Appendix B.2.)
Notably, in the recent work of [29], the authors consider the exact same noise statistic for $H$-consistency in multiclassification—while they operate under structural assumptions of the surrogate and hypothesis class to enable a global convex lower bound, we consider the generic setup with minimal structural constraint, which requires us to introduce the noise condition number in the following.
@@
To provide a somewhat more concrete bound, we optimize over $M$ in Theorem 1, using the crude bound $\kappa(x)\leq 1/\Delta(x)$ on the condition number. By taking $M=1$ for $\mathop{\rm card}(\mathcal{Y})=k=2$ and optimizing $M$ for $k\geq 3$, we may lower bound $\xi_{m,k}$ in the level $(\xi_{m,k},\zeta)$-consistency (Def. 3.2) that in Theorem 1 promises, setting
@@ 3.3 Surrogate consistency examples with majority vote
We collect several examples, of varying levels of concreteness, that allow us to instantiate Theorem 1 and Corollary 3.3. Throughout, we shall assume that $P$ has a noise exponent $\alpha\in[0,1]$, though this is no loss of generality, as Condition (Nα) always holds with $\alpha=0$. We defer proofs for each result in this section to Appendix B.5.
@@ 4 Robustness and consistency for models
The previous section builds off of the now classical theory of surrogate risk consistency, which assumes $\mathcal{F}$ to be the class of all measurable functions. The results there show that aggregation can allow us to “upgrade” consistency so that even if a surrogate $\varphi$ is inconsistent for paired (non-aggregated) data $(X,Y)$, we can achieve level-$(\xi,\zeta)$ consistency (Def. 3.2) with sufficient aggregation. Here, we take a different view of the problem of consistency, considering the consequences of optimizing over a restricted (often parametric) hypothesis class $\mathcal{F}$. Of course, in a well-specified model, obtaining consistency with such a restricted hypothesis class is no issue, but it is unrealistic to assume such a brittle condition. This gives rise to the long-standing challenge of quantifying surrogate consistency when the hypothesis class contains only a subset of the measurable functions [14, 31]. We point out a line of recent work develops $H$-consistency bounds that link excess target risk to surrogate regret within a regret hypothesis class $H$, covering various setups [2, 3, 26, 27, 29]. These results are primarily surrogate-centric. Our contribution is orthogonal, less hinged on the specific class and surrogate, and instead study how label aggregation improves consistency itself. We tackle some of the issues around this, showing that aggregating labels allows consistent estimates in scenarios where consistency might otherwise fail.
@@ 4.1 Consistency failure for binary classification in finite dimensions
Data aggregation methods provide one way to circumvent the inconsistency Proposition 4 highlights. To state the result, define the approximate minimizers
@@ 4.2 Aggregation, consistency, and restricted hypothesis classes
for $\phi$ convex, non-increasing with $\phi^{\prime}(0)<0$, and $\inf_{t}\phi(t)=0$. Zhang [51, Thm. 5] shows that any such loss is consistent over the class $\mathcal{F}=\{f:\mathcal{X}\to\mathbb{R}^{k}\mid\mathbf{1}^{T}f=0\}$. Clearly, the margin-based binary setting in Sec. 4.1 falls into this scenario when we take f(x)=(g(x),g(x))f(x)=(g(x),-g(x)) for a measurable $g$. Additionally, in a parametric setting when $\mathcal{X}=\mathbb{R}^{d}$, if $\mathcal{F}$ consists of linear functions $f(x)=(\langle\theta_{1},x\rangle,\ldots,\langle\theta_{k},x\rangle)$ with $\sum_{i=1}^{k}\theta_{i}=0$, then $\mathcal{F}$ is a (convex) cone.
@@ 4.3 On finite-dimensional multiclass classification
$\varphi(\Theta^{\top}x,y)=-\langle\theta_{y},x\rangle+\log\left({1+\sum_{i=1}^{k-1}\exp(\langle\theta_{i},x\rangle)}\right),$ (15)
@@
Standard results in statistics guarantees both consistency and efficiency when the model is well-specified without aggregation, and when $m\geq 1$ and $k=2$, Cheng et al. [7, Prop. 3] show there exists $t_{m}>0$ such that $\Theta_{m}=t_{m}\Theta^{\star}$ if $X\sim\mathsf{N}(0,I_{d})$ even with a mis-specified link. This implies that in binary classification, even if the link function is incorrect, we can still achieve consistent classification regardless of the aggregation level $m$, as the direction $\Theta^{\star}/\left\|{\Theta^{\star}}\right\|_{2}$ determines consistency. However, as soon as $k\geq 3$ and the true link is slightly mis-specified, risk consistency fails. Fixing a set $\mathcal{T}_{\epsilon}\subset\mathbb{R}^{k-1}$ with Lebesgue measure $\epsilon$, consider
@@
5 Numerical illustrations
In this section, we provide two experiments that corroborate the theory: (i) a non-parametric example where aggregation amplifies surrogate consistency, and (ii) a finite-dimensional multiclass example where aggregation recovers asymptotic consistency under link mis-specification.
Surrogate consistency amplification for truncated quadratic loss.
In binary classification, consider the classical margin-based surrogate $\varphi(f(x),y)=\phi(yf(x))$ where $\phi(\delta)=\max\{1-\delta,0\}^{2}$ is the truncated quadratic loss. By [5, Example 2], the calibration function $\psi$ in Corollary 3.1 is exactly $\psi(\epsilon)=\epsilon^{2}$. Leveraging label aggregation through majority vote with $m$ labels, we can apply Lemma 3.1 taking $\delta=1$. In the regime with no low-noise condition (i.e. $\alpha=0,\mathsf{C}_{\mathsf{MT}}=1$), taking
$\psi_{m}(\epsilon)$ $=\frac{\phi(0)-\phi(\delta)}{16}\left(\epsilon-\sqrt{\frac{32}{m}\log\left({\frac{8(\phi(-\delta)+\phi(0)-\phi(\delta))}{\phi(0)-\phi(\delta)}}\right)}\right)_{+}=\frac{1}{16}\left(\epsilon-\sqrt{\frac{32}{m}\log 40}\right)_{+},$
we have parallel comparison inequalities for non-aggregated data and majority-vote aggregation as
$\psi(R(f)-R^{\star})\leq R_{\varphi}(f)-R_{\varphi}^{\star},\qquad\psi_{m}(R(f)-R^{\star})\leq R_{\varphi,A_{m}}(f)-R_{\varphi,A_{m}}^{\star}.$
Numerically (cf. Fig. 1), the aggregated curve transitions from quadratic toward linear behavior as $m\to\infty$, showing label aggregation upgrades surrogate consistency. Nonetheless, $\psi_{m}$ are generally conservative as they are from the general result in Theorem 1.
Figure 1: Numerical illustrations for truncated quadratic surrogate $\phi(\delta)=\max\{1-\delta,0\}^{2}$, showing comparison inequalities for non-aggregated data $\psi$ (optimal) vs. majority vote aggregation $\psi_{m}$ (suboptimal) when $m\in\{2^{16},2^{20},2^{24},2^{28},2^{32}\}$.
Asymptotic consistency for mis-specified logistic model.
In the second example, we consider logistic regression in Sec. 4.3 when $k=3$ and $d=2$. Let the loss function $\varphi$ be the logistic loss in Eq. (15), with a corrupted logistic link defined as
$\tilde{\sigma}(t)$ $=\sigma^{\textup{lr}}(t){1}\{\left\|{t}\right\|\leq r\}+\sigma^{\textup{lr}}(\alpha t){1}\{\left\|{t}\right\|>r\},$
a piecewise link that matches softmax near the origin and shrinks scores outside a radius $r$. The numerical results are in Fig. 2, showing asymptotic consistency as $m\to\infty$ with synthetic dataset from the corrupted link.
Figure 2: Numerical illustrations for mis-specified logistic models when $d=2$ and $k=3$. The true parameters are $\theta_{1}^{\star}=(0,1)^{\top},\theta_{2}^{\star}=(1,1)^{\top}$. The corrupted link parameters are $r=3$ and $\alpha=0.1$. The plot shows the error of classification rule $\left\|\|T_{m}\|-T^{\star}/\|T^{\star}\|\right\|$ vs. number of labels $m\in\{1,4,16,64,256,1024\}$. For each $m$, the synthetic dataset consist of $n=10,000$ multilabeled data points $(X_{i},(Y_{i1},\cdots,Y_{im}))$ from $\tilde{\sigma}$. We report a $95\%$ error bar on $T=100$ trials for each $n$.
@@ 6 Discussion and future work
Restricted class comparison inequality with label aggregation.
Our paper investigates how label aggregation mitigates regret induced by surrogate losses through comparison inequalities in the non-parametric setting, and establishes statistical consistency in the parametric regime. Motivated by recent developments in $H$-consistency [2, 27, 28, 29], it would be intriguing to extend these techniques to analyze comparison inequalities under restricted hypothesis classes. In particular, [28] shows for smooth convex loss, quadratic growth of $\psi$ is rate optimal—it will be interesting to apply aggregation and upgrade consistency in the context of $\mathcal{H}$-consistency. Such an extension could offer a unified framework linking the advantages of aggregation across parametric and non-parametric settings, while shedding light on how the “size” of the hypothesis class $\mathcal{F}$ influences the effectiveness of aggregation.
@@ A.1 Proof of Proposition 1
Proof  First, we observe that $g_{n}\to g$ pointwise implies that $g_{n}\to g$ uniformly on compacta, and $g$ is convex (see Hiriart-Urruty and Lemaréchal [18, Thm. IV.3.15] and Rockafellar and Wets [38, Thm. 7.17]). This is then equivalent to epigraphical convergence of $g_{n}$ to $g$ [38, Thm. 7.17]. Moreover, as $g_{n}\to g$ uniformly on compacta, if $x_{n}\to x$ then $g_{n}(x_{n})\to g(x)$. Thus, for any $\epsilon_{n}\downarrow 0$, if for a subsequence $n(m)\subset\mathbb{N}$ we have $x_{n(m)}\in\mathop{\epsilon_{n(m)}\textup{-argmin}}g_{n}$, and $x_{n(m)}\to x$, we certainly have $g_{n(m)}(x_{n(m)})\to g(x)$. Consequently [38, Prop. 7.30] we have
@@
That the limit supremum is non-empty is then a consequence of [38, Thm. 7.33], as the convex functions $g_{n}$ must be coercive as they are convex and $g$ is. ∎
@@ B.2 The equivalence of the Mammen-Tsybakov conditions (Nα) and (Mβ)
We show the analogue of Bartlett et al. [5, Thm. 3], essentially mimicking their proof, but including it for completeness.
@@ B.3 Proof of Theorem 1
Consider the second term in the bound (21). For any convex function $0\leq\psi$ with $\psi(0)=0$, $\epsilon\mapsto\psi(\epsilon)/\epsilon$ is non-decreasing on $\epsilon>0$ (cf. [18, Ch. 1]). This implies
@@ B.5 Proofs for the identifiable surrogate losses
That $\phi(\delta)<0$ is immediate because $\phi$ is non-increasing by assumption, and the monotonicity properties of convex functions [18, Ch. 1] guarantee it strictly decreases near 0. For $y\in\{\pm 1\}$, we take $s_{y}=\delta y$ and $a_{y}=y$, and
@@ D.2 Proof of Theorem 2
the convex conjugate of the gap functional on well-separated distributions. Then Zhang [51, Prop. 25] shows that $\overline{\psi}_{\varphi}^{**}(\epsilon)>0$ whenever $\overline{\psi}_{\varphi}(\epsilon)>0$.